OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). -21 -3 3 21

OpenStudy (ciarán95):

The simplest method of working this out is to first find g(1), that is substitute in '1' for everywhere you find 'x' in g(x) and evaluate. Now, as g(1) will just work out to be a single number, then finding f(g(1)) should be straight forward. Simply substitute in your answer for 'g(1)' everywhere you find 'x' in f(x) = x - 2. Hope that helps! :)

OpenStudy (anonymous):

g(-1)=-1 f(g(-1))=-1-2=-3 theres Your answer. :)

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33

OpenStudy (anonymous):

Thats Why maths is so easy. The way I did was the concept and if you you do anything conceptually you will never be wrong.

OpenStudy (anonymous):

Well, I tried to plug it in but I was not sure if it was correct the way I plugged it in.

OpenStudy (anonymous):

You were right BTW. Have confidence in what you do if the result is right.